# Turing Video-based Cognitive Tests to Handle Entangled Concepts

### v1

@@ Abstract
We have proved in both human-based and computer-based tests that natural concepts generally ‘entangle’ when they combine to form complex sentences, violating the rules of classical compositional semantics. In this article, we present the results of an innovative video-based cognitive test on a specific conceptual combination, which significantly violates the Clauser–Horne–Shimony–Holt version of Bell’s inequalities (‘CHSH inequality’). We also show that collected data can be faithfully modelled within a quantum-theoretic framework elaborated by ourselves and a ‘strong form of entanglement’ occurs between the component concepts. While the video-based test confirms previous empirical results on entanglement in human cognition, our ground-breaking empirical approach surpasses language barriers and eliminates the need for prior knowledge, enabling universal accessibility. Finally, this transformative methodology allows one to unravel the underlying connections that drive our perception of reality. As a matter of fact, we provide a novel explanation for the appearance of entanglement in both physics and cognitive realms.
Keywords: human cognition, concept combinations, Bell’s inequalities, quantum modelling, entanglement
@@ 1 Introduction
Completely Automated Public Turing test to tell Computers and Humans Apart (CAPTCHA [^4] The term “CAPTCHA” was coined in 2000 by Luis von Ahn, Manuel Blum, Nicholas Hopper and John Langford of Carnegie Mellon University, see, e.g., the webpage http://www.captcha.net.) have been widely used by websites to safeguard their services from bot-driven activities. Relatively simple tasks for people, but challenging for computers, are presented in CAPTCHA, such as clicking in a designated area, recognizing letters or numbers that are stretched, selecting objects in an image. Their simple yet effective mechanisms have contributed to maintain poll integrity, control registration, prevent ticket inflation, and counter false comments, ultimately ensuring a more secure and trustworthy online experience for users.
Image-based CAPTCHAs require both image recognition and semantic classification, making them harder for bots to understand than text-based one. In addition, by collecting data from CAPTCHAs, one could train machine learning models.
@@
It is commonly accepted that human perception and thought are essentially synthetic processes. We immediately form a ‘Gestalt’, or a general idea, of the object. Gestaltic patterns seem to be the primary basis of rational activity as well. Gestalt-thinking is not well captured by the analytical and compositional framework of classical logical semantics, which essentially deduces the meaning of a composite expression from the meanings of its component parts (‘principle of compositionality’, see, e.g., [1]). For an accurate study of natural languages and creative contexts, where holistic and ambiguous characteristics seem to be significant, classical semantics is no longer very useful. One example in this regard is the last line of Giacomo Leopardi’s poem L’Infinito, “E ’l naufragar m’è dolce in questo mare” (And drowning in this sea is sweet to me, i.e. I think it’s delightful to dedicate myself to the contemplation and meditation of the infinite). The meanings of the component terms “naufragar” (drowning), “dolce” (sweet), and “mare” (sea) do not correlate to the meanings they have in this context, which appears to be the fundamental cause of the poetic outcome in this instance. By the way, the poem refers to the settlement of Recanati, which is not near a sea. Nonetheless, our words’ common meanings continue to be used and are loosely connected to the metaphorical meanings that the poem as a whole evokes, a semantic conundrum that frequently occurs in poetry and musical compositions, where meanings are inherently holistic, contextual and ambiguous.
@@
Similar issues occur whenever people combine individual concepts to form conceptual combinations or more complex linguistic expressions as sentences and texts. A large empirical literature indeed reveals that the principle of compositionality is systematically violated and, more generally, concepts exhibit aspects of ‘inherent vagueness’, ‘contextuality’ and ‘emergence’ which prevent them to be modelled within classical Boolean logical and classical Kolmogorovian probabilistic structures (see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10] and references therein).
@@
Recently, various approaches have been put forward to theoretically cope with this structural inability of classical compositional semantics to handle the dynamics of human concepts. Remarkably, some of these approaches use the logico-mathematical formalism of quantum theory, detached from its physical interpretation, as a modelling tool to represent conceptual meaning. This research fits a growing research programme that applies quantum structures in the mathematical modelling of cognitive processes, with relevant extensions to information retrieval processes (see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and references therein).
@@
In particular, in the ‘holistic quantum computational semantics’, meanings are typically represented by quantum superpositions of meanings and are treated as fundamentally dynamical objects. Any global meaning determines partial meanings, which are frequently vaguer than the global one. The majority of research on quantum computational logics has focused on sentential logics, whose alphabet is made up of atomic sentences and logical connectives. A first-order epistemic quantum computational logic with a semantic characterization that can express sentences like “the animal acts,” “the animal eats the food”, etc. was presented in [26].
An alternative but similar approach considers a concept as an entity whose meaning is incorporated into a given state, whose state can change under the influence of a context, and represents conceptual entities in the formalism of quantum theory in Hilbert space [27].
Within the two quantum approaches above, empirical and theoretical studies have been carried out with the aim of identifying ‘quantum entanglement’ in the combination of natural concepts [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. In physics, entanglement is a statistical property of a composite system, or ‘entity’, made up, in the simplest case, of two individual entities. Entanglement is typically detected through the violation of suitable inequalities, called ‘Bell’s inequalities’, the ‘Clauser–Horne–Shimony–Holt (CHSH) inequality’ [40, 41] being one of these inequalities. More precisely, we have recently performed various tests, including text-based cognitive tests on human participants, information retrieval tests on corpora of documents, and image retrieval tests on web search engines, which confirm that the CHSH inequality is systematically violated when two concepts combine [30, 31, 33, 34, 35, 36, 37, 38, 39].
@@
In this article, we deepen and widely extend the research above and present the results of a video-based cognitive test that we have recently performed on the conceptual combination The Animal Acts, considered as a combination of the concepts Animal and Acts, [^5] In this article, we write concepts using capital letters and italics, e.g., Animal, Fruit, Vegetables, etc. This is frequently the way of referring to concepts in cognitive psychology. where the term “acts” refers to one of the sounds that can be made by an animal. The test used the videos produced by CAPTCHA through artificial intelligence (AI) and showed to the participants, who had to judge among the videos the best examples of The Animal Acts. As such, the test is more realistic and effective than the previous text-based tests, as it surpasses language barriers and eliminates the need for prior knowledge, thus enabling universal accessibility. We show that the CHSH inequality is significantly violated in the test, which reveals the presence of entanglement between the component concepts. We also work out a quantum theoretical model in Hilbert space for the statistical data (‘judgement probabilities’), which reveals that entanglement occurs at both state and measurement levels, hence it is even stronger than the entanglement that is typically detected in quantum physics tests.
@@
The presence of entanglement can be naturally explained if one observes that both concepts Animal and Acts carry meaning, but also the combination The Animal Acts carries its own meaning, and this meaning is not simply related to the separate meanings of Animal and Acts as prescribed by a classical compositional semantics. It also contains emergent meaning almost completely caused by its interaction with the wide overall context. We have called the complex mechanism by which meaning is attributed to the combined concept ‘contextual updating’, and it occurs at the level of entanglement creation [42].
@@
In Section 2, we illustrate the general setting of a Bell-type test for the empirical detection of entanglement in both physical and conceptual domains. The violation of the CHSH inequality is generally considered as conclusive towards the presence of entanglement in the given situation. We however observe that situations exist in which entanglement cannot only be attributed to the state of a composite entity, but also the measurements need to be considered to be entangled. This typically occurs in conceptual domains.
@@
In Section 4, we work out a quantum mathematical representation in Hilbert space for the data collected in Section 3 and show that the violation of the CHSH inequality can be explained in terms of a strong form of entanglement involving both states and measurements.
In Section 5, we finally provide a theoretical analysis of the results obtained in the test and the ensuing modelling, and establish some deep connections between the non-classical mechanism of meaning attribution and the appearance of entanglement in conceptual domains.
@@ 2 Identifying entanglement in physics and cognition
We present in this section the typical way in which entanglement is empirically detected in physical and cognitive domains.
In physics, entanglement is a statistical property of a composite entity, which can be identified by means of the empirical violation of Bell’s inequalities [40, 41, 43]. One of Bell’s inequalities that is more suited for empirical control is the CHSH inequality [41]. The violation of Bell’s inequalities is generally interpreted by saying that, due to entanglement, micro-physical entities exhibit genuinely non-classical aspects, as ‘nonseparability’ and ‘contextuality’, where the latter also occurs when the component entities are far apart in space, an aspect known as ‘nonlocality’. In addition, entanglement produces statistical correlations between the component entities that cannot be reproduced by any classical model of probability satisfying the axioms of Kolmogorov [44, 45]. Because of its deep physical implications, entanglement has been considered as the distinctive trait of quantum theory since its early days [46].
The empirical setting for the identification of entanglement refers to a ‘Bell-type test’, since the seminal papers by Einstein–Podolsky–Rosen (EPR), David Bohm and John Bell [40, 41, 43, 47, 48]. Let us consider a composite physical entity ${S}_{12}$, prepared in an initial state $p$, and such that the individual entities ${S}_{1}$ and ${S}_{2}$ can be recognized as component parts of ${S}_{12}$. Next, let us perform the coincidence measurements $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ on $S_{12}$, where the coincidence measurement $XY$ consists in performing the measurement $X$ on $S_{1}$, with possible outcomes $X_{1}$ and $X_{2}$, and the measurement $Y$ on $S_{2}$, with possible outcomes $Y_{1}$ and $Y_{2}$, with $X=A,A^{\prime}$, $Y=B,B^{\prime}$. The component entities $S_{1}$ and $S_{2}$ have interacted in the past, but are spatially separated when the coincidence measurements are performed. If $X_{1},X_{2},Y_{1},Y_{2}$ can only be ±$\pm 1$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, the expected values of $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ are the correlation functions $E(A,B)$, $E(A,B^{\prime})$, $E(A^{\prime},B)$ and $E(A^{\prime},B^{\prime})$, respectively. One can then prove that, under the reasonable, in classical physics, assumption of ‘local separability’, or ‘local realism’ [47], the CHSH inequality, namely,
@@
should be satisfied [41]. We call ‘CHSH factor’ the quantity
@@
According to modern manuals of quantum theory, the individual entities $S_{1}$ and $S_{2}$ are associated with the complex Hilbert spaces ${\mathscr{H}}_{1}$ and ${\mathscr{H}}_{2}$, respectively. In this case, both ${\mathscr{H}}_{1}$ and ${\mathscr{H}}_{2}$ are isomorphic to the complex Hilbert space $\mathbb{C}^{2}$ of all ordered couples of complex numbers. The composite entity $S_{12}$ is instead associated with the tensor product Hilbert space ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$. In this case, ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ is isomorphic to the complex Hilbert space $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. The possible (pure) states of $S_{1}$ and $S_{2}$ are represented by unit vectors of ${\mathscr{H}}_{1}$ and ${\mathscr{H}}_{2}$, respectively, and the measurements that can be performed on $S_{1}$ and $S_{2}$ are represented by self-adjoint operators on ${\mathscr{H}_{1}}$ and ${\mathscr{H}}_{2}$, respectively. However, ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ also contains vectors that cannot be written as the tensor product of a unit vector of ${\mathscr{H}_{1}}$ and a unit vector of ${\mathscr{H}_{2}}$. These non-product vectors of ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ are said to represent ‘non-product’, or ‘entangled’, states of $S_{12}$. Analogously, the self-adjoint operators of ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ are not limited to operators that are the tensor product of a self-adjoint operator of ${\mathscr{H}}_{1}$ and a self-adjoint operator of ${\mathscr{H}}_{2}$. In these cases, at least one eigenvector of these non-product self-adjoint operators represents an entangled state. These non-product self-adjoint operators are said to represent ‘non-product’, or ‘entangled’, measurements of $S_{12}$ [43].
The CHSH inequality in Equation (1) is manifestly violated in quantum theory and, when a violation occurs, it is due to the presence of entanglement between the component entities $S_{1}$ and $S_{2}$ that can be recognized as component parts of $S_{12}$. Equivalently, in the absence of entanglement, the inequality in Equation (1) would not be violated. The typical situation in which the inequality in Equation (1) is violated in quantum theory consists in the state $p$ of the composite entity $S_{12}$ being the ‘singlet spin state’, which is an example of a ‘maximally entangled state’ [43], and the coincidence measurements $XY$ being product measurements, $X=A,A^{\prime}$, $B^{\prime}B^{\prime}$. This situation entails a CHSH factor equal to $\Delta_{QMC}=2\sqrt{2}\approx 2.83$, known as ‘Cirel’son’s bound’ [49, 50]. The situation, however, also requires a special ‘symmetry requirement on joint probabilities’, namely, the ‘marginal law of Kolmogorovian probability being satisfied’, i.e. for every $i,j=1,2$,
@@
In Equation (3), $\mu(X_{i}Y_{j})$ ($\mu(X_{i}Y^{\prime}_{j})$) is the joint probability of obtaining the outcomes $X_{i}$ in a measurement of $X$ on $S_{1}$ and $Y_{j}$ ($Y^{\prime}_{j}$) in a measurement of $Y$ ($Y^{\prime}$) on $S_{2}$, $X=A,A^{\prime}$, $Y,Y^{\prime}=B,B^{\prime}$, $Y^{\prime}\neq Y$. In Equation (4), $\mu(X_{i}Y_{j})$ ($\mu(X^{\prime}_{i}Y_{j})$) is the joint probability of obtaining the outcomes $X_{i}$ ($X^{\prime}_{i}$) in a measurement of $X$ ($X^{\prime}$) on $S_{1}$ and $Y_{j}$ in a measurement of $Y$ ($Y^{\prime}$) on $S_{2}$, $X,X^{\prime}=A,A^{\prime}$ and $Y=B,B^{\prime}$, $X^{\prime}\neq X$.
@@
The needs of quantum computation and quantum information have intensified the theoretical research on Bell’s inequalities (see, e.g., [43, 51]) as well as the research on the empirical consequences of entanglement (see, e.g., [51, 53, 54]), confirming the predictions of quantum theory.
@@
In the meanwhile, it has become evident that the genuinely quantum aspects of contextuality, entanglement, indistinguishability, interference, and superposition, are not peculiar of micro-physical entities, as they also occur in the quantum mathematical modelling of complex cognitive processes, e.g., the processes involving judgement, decision, perception and language (see Section 1). In particular, both experimental and theoretical research has been dedicated to the identification of entanglement in the combination of natural concepts (see again Section 1), and our research team has provided substantial contributions along both lines of this investigation, as follows.

### v2

# Turing Video-based Cognitive to Handle Tests Entangled Concepts

@@ Abstract
We have proved in both human-based and computer-based tests that natural concepts generally ‘entangle’ when they combine to form complex sentences, violating the rules of classical compositional semantics. In this article, we present the results of an innovative video-based cognitive test on a specific conceptual combination, which significantly violates the Clauser–Horne–Shimony–Holt version of Bell’s inequalities (‘CHSH inequality’). We also show that collected data can be faithfully modelled within a quantum-theoretic framework elaborated by ourselves and a ‘strong form of entanglement’ occurs between the component concepts. While the video-based test confirms previous empirical results on entanglement in human cognition, our empirical approach surpasses language barriers and eliminates the need for prior knowledge, enabling universal accessibility. Finally, this transformative methodology allows one to unravel the underlying connections that drive our perception of reality. As a matter of fact, we provide a novel explanation for the appearance of entanglement in both physics and cognitive realms.
Keywords: human cognition, concept combinations, entanglement, Bell’s inequalities, quantum modelling
@@ 1 Introduction
Completely Automated Public Turing test to tell Computers and Humans Apart (CAPTCHA [^4] The term “CAPTCHA” was coined in 2000 by Luis von Ahn, Manuel Blum, Nicholas Hopper and John Langford, researchers at Carnegie Mellon University. Additional information can be found, e.g., on the webpage http://www.captcha.net.) have been widely used by websites to safeguard their services from bot-driven activities. Relatively simple tasks for people, but challenging for computers, are presented in CAPTCHA, such as clicking in a designated area, recognizing letters or numbers that are stretched, selecting objects in an image. Their simple yet effective mechanisms have contributed to maintain poll integrity, control registration, prevent ticket inflation, and counter false comments, ultimately ensuring a more secure and trustworthy online experience for users. Image-based CAPTCHAs require both image recognition and semantic classification, making them harder for bots to understand than text-based one. In addition, by collecting data from CAPTCHAs, one could train machine learning models.
@@
It is commonly accepted that human perception and thought are essentially synthetic processes. We immediately form a ‘Gestalt’, or a general idea, of the object. Gestaltic patterns seem to be the primary basis of rational activity as well. Gestalt-thinking is not well captured by the analytical and compositional framework of classical logical semantics, which essentially deduces the meaning of a composite expression from the meanings of its component parts (‘principle of compositionality’, see, e.g., Carnap, (1947)). For an accurate study of natural languages and creative contexts, where holistic and ambiguous characteristics seem to be significant, classical semantics is no longer very useful. One example in this regard is the last line of Giacomo Leopardi’s poem ‘L’Infinito’, “E ’l naufragar m’è dolce in questo mare” (And drowning in this sea is sweet to me, i.e. I think it’s delightful to dedicate myself to the contemplation and meditation of the infinite). The meanings of the component terms “naufragar” (drowning), “dolce” (sweet), and “mare” (sea) do not correlate to the meanings they have in this context, which appears to be the fundamental cause of the poetic outcome in this instance. By the way, the poem refers to the settlement of Recanati, which is not near a sea. Nonetheless, our words’ common meanings continue to be used and are loosely connected to the metaphorical meanings that the poem as a whole evokes, a semantic conundrum that frequently occurs in poetry and musical compositions, where meanings are inherently holistic, contextual and ambiguous.
@@
Similar issues occur whenever people combine individual concepts to form conceptual combinations or more complex linguistic expressions as sentences and texts. A large empirical literature indeed reveals that the principle of compositionality is systematically violated and, more generally, concepts exhibit aspects of ‘inherent vagueness’, ‘contextuality’ and ‘emergence’ which prevent them from being modelled within classical Boolean logical and classical Kolmogorovian probabilistic structures (see, e.g., Aerts, 2009a ; Aerts, 2009b ; Coecke Sadrzadeh & Clark, (2010); Busemeyer & Bruza, (2012); Aerts et al., (2013); Aerts Gabora & Sozzo, (2013); Bruza et al., (2016); Dalla Chiara et al., 2015a and references therein).
@@
Recently, various approaches have been put forward to theoretically cope with this structural inability of classical compositional semantics to handle the dynamics of human concepts. Remarkably, some of these approaches use the logico-mathematical formalism of quantum theory, detached from its physical interpretation, as a modelling tool to represent conceptual meaning. This research fits a growing programme that applies quantum structures in the mathematical modelling of cognitive processes, with relevant extensions to information retrieval processes (see, e.g., Aerts, 2009a ; Aerts, 2009b ; Coecke Sadrzadeh & Clark, (2010); Busemeyer & Bruza, (2012); Aerts et al., (2013); Aerts Gabora & Sozzo, (2013); Bruza et al., (2016); Dalla Chiara et al., 2015a ; Pothos & Busemeyer, (2009); Frommholz et al., (2010); Khrennikov, (2010); Piwowarski et al., (2010); Buccio Melucci & Song, (2011); Zellhöfer et al., (2011); Haven & Khrennikov, (2013); Dalla Chiara et al., 2015c ; Kvam et al., (2015); Melucci, (2015); Aerts et al., (2017, 2018); Aerts Haven & Sozzo, (2018); Pisano & Sozzo, (2020); Aerts et al., (2021) and references therein).
@@
In particular, in the ‘holistic quantum computational semantics’ approach, meanings are typically represented by quantum superpositions of meanings and are treated as fundamentally dynamical objects. Any global meaning determines partial meanings, which are frequently vaguer than the global one. The majority of research on quantum computational logics has focused on sentential logics, whose alphabet is made up of atomic sentences and logical connectives. A first-order epistemic quantum computational logic with a semantic characterization that can express sentences like “the animal acts,” “the animal eats the food”, etc. was presented in Dalla Chiara et al., 2015b . An alternative, but similar, approach considers a concept as an entity whose meaning is incorporated into a given state, whose state can change under the influence of a context, and represents conceptual entities in the formalism of quantum theory in Hilbert space (Aerts Sassoli de Bianchi & Sozzo,, 2016). Within these two quantum approaches, empirical and theoretical studies have been carried out with the aim of identifying ‘quantum entanglement’ in the combination of natural concepts (Bruza et al., 2009a, ; Bruza et al., 2009b, ; Aerts & Sozzo,, 2011, 2014; Gronchi & Strambini,, 2017; Aerts Arguëlles,, 2018; Aerts et al.,, 2019; Beltran & Geriente,, 2019; Aerts Arguëlles & Sozzo,, 2020; Aerts et al.,, 2021; Aerts et al., 2023b, ; Bertini Lemporini & Moriani,, 2023).
In physics, the notion of entanglement was introduced by Erwin Schrödinger in a letter to Albert Einstein, after reading the 1935 article where Einstein, with his two collaborators Boris Podolsky and Nathan Rosen, analysed a specific type of correlations, nowadays known as ‘EPR correlations’, appearing in a composite quantum system, or ‘entity’, made up of two individual entities (Einstein Podolsky & Rosen,, 1935). In the same year, Schrödinger published an article on the matter, where he officially introduced the notion of entanglement as a phenomenon that, while appearing for two quantum entities separated in space, seems to indicate that the entities are ‘non-separated’ (Schrödinger,, 1935). This made Einstein use the expression ‘spooky action at-a-distance’ when talking about the phenomenon. In 1951, David Bohm introduced the archetypical situation of two spin 1/2 quantum entities in the singlet spin state of total spin equal to zero and flying apart (Bohm,, 1951). If one imagines that, (i) if one spin is forced “up” by the measurement apparatus applied to it, then, as a consequence, the other spin is ‘immediately’ forced “down”, even when not measured upon, and (ii) this ‘mechanism’ keeps occurring for two quantum entities independently of their distance, the phenomenon could be called indeed a spooky action at-a-distance. It is John Bell who has the great merit of making the question open to experimental verification by deriving an inequality that should not be violated if certain empirical tests are performed on EPR correlated quantum systems while quantum theory predicts that the inequalities derived by Bell will be violated (Bell,, 1964). This has also had a relevance on the so-called ‘hidden variables programme’, initiated after the EPR article, because these results entail that any hidden variable completion of quantum theory has unavoidably to be non-local.
After the seminal work of Bell, several similar inequalities have been derived, and one typically refers to them as ‘Bell’s inequalities’ (see, e.g., Brunner et al., (2014)), one of them being the ‘Clauser–Horne–Shimony–Holt (CHSH) inequality’ (Clauser et al.,, 1969). The CHSH inequality is particularly interesting, as it is well suited, not only for empirical tests of non-locality in quantum physics, but also as a test to detect the presence of entanglement, within and beyond physical domains. Indeed, in the absence of entanglement, the CHSH inequality would not be violated. [^5] We notice that the violation of Bell’s inequalities, while being a sufficient condition for the detection of entanglement, is not necessary because, for a given Bell’s inequality, entangled states exist which do not violate it. As a matter of fact, we have performed in the last decade various tests, including text-based cognitive tests on human participants, information retrieval tests on corpora of documents, and image retrieval tests on web search engines, which confirm that the CHSH inequality is systematically violated in cognitive domains, particularly, when two concepts combine (Aerts & Sozzo,, 2011, 2014; Aerts Arguëlles,, 2018; Aerts et al.,, 2019; Beltran & Geriente,, 2019; Aerts Arguëlles & Sozzo,, 2020; Aerts et al.,, 2021; Aerts et al., 2023b, ; Bertini Lemporini & Moriani,, 2023).
@@
In this article, we deepen and extend the investigation above and present the results of a video-based cognitive test that we have recently performed on the conceptual combination The Animal Acts, considered as a combination of the concepts Animal and Acts, [^6] In this article, we write concepts using capital letters and italics, e.g., Animal, Fruit, Vegetables, etc. This is frequently the way of referring to concepts in cognitive psychology. where the term “acts” refers to one of the sounds that can be made by an animal. The test used the videos produced by CAPTCHA through artificial intelligence (AI) and showed to the participants, who had to judge among the videos the best examples of The Animal Acts. As such, the test is more realistic and effective than the previous text-based tests, as it surpasses language barriers and eliminates the need for prior knowledge, thus enabling universal accessibility. We show that the CHSH inequality is significantly violated in the test, which reveals the presence of entanglement between the component concepts. We also work out a quantum theoretical model in Hilbert space for the statistical data (‘judgement probabilities’), which reveals that entanglement occurs at both state and measurement levels, hence it is even stronger than the entanglement that is typically detected in quantum physics tests.
@@
The presence of entanglement can be naturally explained if one observes that both concepts Animal and Acts carry meaning, but also the combination The Animal Acts carries its own meaning, and this meaning is not simply related to the separate meanings of Animal and Acts as prescribed by a classical compositional semantics. It also contains emergent meaning almost completely caused by its interaction with the wide overall context. We have called the complex mechanism by which meaning is attributed to the combined concept ‘contextual updating’, and it occurs at the level of entanglement creation (Aerts et al., 2023a, ).
@@
In Section 2, we illustrate the general setting of a Bell-type test for the empirical detection of entanglement in both physical and cognitive, i.e. conceptual and linguistic, domains. The violation of the CHSH inequality is generally considered as conclusive towards the presence of entanglement in the given situation. We however observe that situations exist in which entanglement cannot only be attributed to the state of a composite entity, but also the measurements need to be considered to be entangled. This typically occurs in cognitive domains.
In Section 3, we explain the general reasons which led us to put forward a theoretical approach that diverges in some aspects from more known approaches.
@@
In Section 5, we work out a quantum mathematical representation in Hilbert space for the data collected in Section 4 and show that the violation of the CHSH inequality can be attributed to the presence of a strong form of entanglement involving both states and measurements, as anticipated in Section 2.
In Section 6, we finally provide a theoretical analysis of the results obtained in the test and the ensuing modelling, and establish some deep connections between the non-classical mechanism of meaning attribution and the appearance of entanglement in cognitive domains.
@@ 2 Identifying entanglement in physics and cognition
We present in this section the typical way in which entanglement is theoretically understood and empirically detected in physical and cognitive, i.e. conceptual and linguistic, domains.
In Section 1, we have introduced entanglement as a statistical property of a composite entity, which can be identified by means of the empirical violation of Bell’s inequalities (Bell,, 1964; Brunner et al.,, 2014; Clauser et al.,, 1969). We have also seen that an inequality that is particularly suited for empirical control is the CHSH inequality (Clauser et al.,, 1969). We remind that the violation of Bell’s inequalities is typically interpreted by saying that, due to entanglement, micro-physical entities exhibit genuinely non-classical aspects, as non-separability, contextuality and non-locality. We add that the statistical correlations produced in the presence of entanglement cannot be reproduced by any classical model of probability satisfying the axioms of Kolmogorov (Accardi & Fedullo,, 1982; Pitowsky,, 1989). Because of its deep physical implications, entanglement has been considered as the distinctive trait of quantum theory since its early days (Schrödinger,, 1935).
The empirical setting for the identification of entanglement refers to a ‘Bell-type test’ (Einstein Podolsky & Rosen,, 1935; Bohm,, 1951; Bell,, 1964; Brunner et al.,, 2014; Clauser et al.,, 1969), which can be outlined as follows. Let us consider a composite physical entity ${S}_{12}$, prepared in an initial state $p$, and such that the component entities ${S}_{1}$ and ${S}_{2}$ can be recognized as component parts of ${S}_{12}$. Next, let us perform the joint measurements $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ on $S_{12}$, where the joint measurement $XY$ consists in performing the measurement $X$ on $S_{1}$, with possible outcomes $X_{1}$ and $X_{2}$, and the measurement $Y$ on $S_{2}$, with possible outcomes $Y_{1}$ and $Y_{2}$, with $X=A,A^{\prime}$, $Y=B,B^{\prime}$. The component entities $S_{1}$ and $S_{2}$ have interacted in the past, but are spatially separated when the joint measurements are performed. If $X_{1},X_{2},Y_{1},Y_{2}$ can only be ±$\pm 1$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, the expected values of $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ are the correlation functions $E(A,B)$, $E(A,B^{\prime})$, $E(A^{\prime},B)$ and $E(A^{\prime},B^{\prime})$, respectively. One can then prove that, under the reasonable, in classical physics, assumption of ‘local separability’, or ‘local realism’ (Einstein Podolsky & Rosen,, 1935), the CHSH inequality, namely,
@@
should be satisfied (Clauser et al.,, 1969). We call ‘CHSH factor’ the quantity
@@
According to modern manuals of quantum theory, the component entities $S_{1}$ and $S_{2}$ are associated with the complex Hilbert spaces ${\mathscr{H}}_{1}$ and ${\mathscr{H}}_{2}$, respectively. In this case, both ${\mathscr{H}}_{1}$ and ${\mathscr{H}}_{2}$ are isomorphic to the complex Hilbert space $\mathbb{C}^{2}$ of all ordered pairs of complex numbers. The composite entity $S_{12}$ is instead associated with the tensor product Hilbert space ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$. In this case, ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ is isomorphic to the complex Hilbert space $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. The possible (pure) states of $S_{1}$ and $S_{2}$ are represented by unit vectors of ${\mathscr{H}}_{1}$ and ${\mathscr{H}}_{2}$, respectively, and the measurements that can be performed on $S_{1}$ and $S_{2}$ are represented by self-adjoint operators on ${\mathscr{H}_{1}}$ and ${\mathscr{H}}_{2}$, respectively. However, ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ also contains vectors that cannot be written as the tensor product of a unit vector of ${\mathscr{H}_{1}}$ and a unit vector of ${\mathscr{H}_{2}}$. These non-product vectors of ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ are said to represent ‘non-product’, or ‘entangled’, states of $S_{12}$. Analogously, the self-adjoint operators of ${\mathscr{H}_{1}}\otimes{\mathscr{H}}_{2}$ are not limited to operators that are the tensor product of a self-adjoint operator of ${\mathscr{H}}_{1}$ and a self-adjoint operator of ${\mathscr{H}}_{2}$. In these cases, at least one eigenvector of these non-product self-adjoint operators represents an entangled state. These non-product self-adjoint operators are said to represent ‘non-product’, or ‘entangled’, measurements of $S_{12}$ (Brunner et al.,, 2014).
The CHSH inequality in Equation (1) is manifestly violated in quantum theory and, when a violation occurs, it is due to the presence of entanglement between the component entities $S_{1}$ and $S_{2}$ that can be recognized as component parts of $S_{12}$. Equivalently, in the absence of entanglement, the inequality in Equation (1) would not be violated. As we have anticipated in Section 1, the typical situation in which the inequality in Equation (1) is violated in quantum theory consists in the state $p$ of the composite entity $S_{12}$ being the singlet spin state (an example of a maximally entangled state (Brunner et al.,, 2014)) and the joint measurements $XY$ being product measurements, $X=A,A^{\prime}$, $B^{\prime}B^{\prime}$. This situation entails a CHSH factor equal to $\Delta_{QMC}=2\sqrt{2}\approx 2.83$, known as ‘Cirel’son’s bound’ (Cirelson,, 1980, 1993). A situation involving product measurements automatically satisfies the ‘marginal law of Kolmogorovian probability’, i.e. the conditions that, for every $i,j=1,2$,
@@
In Equation (3), $\mu(X_{i}Y_{j})$ ($\mu(X_{i}Y^{\prime}_{j})$) is the joint probability of obtaining the outcomes $X_{i}$ in a measurement of $X$ on $S_{1}$ and $Y_{j}$ ($Y^{\prime}_{j}$) in a measurement of $Y$ ($Y^{\prime}$) on $S_{2}$, $X=A,A^{\prime}$, $Y,Y^{\prime}=B,B^{\prime}$, $Y^{\prime}\neq Y$. In Equation (4), $\mu(X_{i}Y_{j})$ ($\mu(X^{\prime}_{i}Y_{j})$) is the joint probability of obtaining the outcomes $X_{i}$ ($X^{\prime}_{i}$) in a measurement of $X$ ($X^{\prime}$) on $S_{1}$ and $Y_{j}$ in a measurement of $Y$ ($Y^{\prime}$) on $S_{2}$, $X,X^{\prime}=A,A^{\prime}$ and $Y=B,B^{\prime}$, $X^{\prime}\neq X$. The conditions in Equations (3) and (4) are also known as ‘no-signaling’, or ‘marginal selectivity’, conditions, and have been widely studied in both physical (see, e.g., Brunner et al., (2014); Peres & Terno, (2004); Horodecki Horodecki & Horodecki, (2009)) and cognitive literature (see, e.g., de Barros, (2015); Dzhafarov & Kujala, (2016)). A violation of these conditions has been traditionally considered as problematical by physicists, as it would open up the possibility of a faster-than-light communication. We will come back to this point in Section 3.
@@
The needs of quantum computation and quantum information have intensified the theoretical research on Bell’s inequalities (see, e.g., Brunner et al., (2014); Genovese, (2005)) as well as the research on the empirical consequences of entanglement (see, e.g., Genovese, (2005); Giustina et al., (2013); Christensen et al., (2013)), confirming the predictions of quantum theory.
@@
In the meanwhile, it has become evident that the genuinely quantum aspects of contextuality, entanglement, indistinguishability, interference, and superposition, are not peculiar of micro-physical entities, as they also occur in the quantum mathematical modelling of complex cognitive processes, e.g., the processes involving categorization, judgement, decision, perception, and language. In particular, both experimental and theoretical research has been dedicated to the identification of entanglement in the combination of natural concepts, and our research team has provided substantial contributions along both lines of this investigation, as follows.